Someone get 750 naira a day, does o/time for 3hrs @ 1.5hrs calculation. The result pls Thanks David Bassey.
Answer by Theo(13342)

he makes 750 naira a day.
he works overtime for 3 hours at 1.5 time the regular pay.
this all depends on how many hours his regular day is.

if his regular day is 8 hours, then 3/8 * 750 * 1.5 = 421.075 naira.

consider that 750/8 = 93.75 naira per regular hour * 1.5 = 140.625 naira per overtime hour.

3 time 140.625 = 421.875 naira.

if his regular day is 12 hours, then 3/12 * 750 * 1.5 = 281.25.

consider that 750 /12 = 62.5 naira per regular hour * 1.5 = 93.75 naira per overtime hour.

3 times 93.75 = 281.25.

you can't really answer this until you know what the regular pay per hour is.

the general equation would be 750 divided by the number of regular hours in the day * 3 * 1.5